a store mixes kenyan coffee worth $10 per kilogram and turkish coffee worth $13 per kg. the mixture is to sell for $12 per kg. how much to make a 600kg mixture?
Answer by mananth(16946)

Price Price ---------------- quantity
Kenyan coffee $10 ----------------x kg
Turkish coffee $13 ----------------600-x kg
Mixture $ 12.00 ---------------- 600

10x+13( 600-x)= 12*600
10x+7800-13x=7200
10x-13x =7200-7800
-3x=-600
/-3
x=200kg $10.00 Kenyan coffee
Balance 400 kg 13.00% Turkish coffee
m.ananth@hotmail.ca